OpenStudy (anonymous):

Solve for x. 4(x – b) = x i don't get how to do it i have tried i just dont understand

OpenStudy (anonymous):

should the answer be in terms of b

OpenStudy (anonymous):

yes

OpenStudy (phi):

First, use the distributive property: 4(a+b) is the same as 4*a+4*b For example, 4(2+3)= 4*2+ 4*3= 8+12= 20 or 4(5)= 20 in your case: 4(x-b)=x you get: 4x-4b=x

OpenStudy (phi):

can you solve 4x-4b= x ?? subtract x from both sides of the equation add 4b to both sides of the equation. If this does not make sense, watch this http://www.khanacademy.org/video/simple-equations?playlist=Algebra and the next 2 after it.

OpenStudy (anonymous):

thanks alot